OpenStudy (anonymous):

cot 91* cot92* cot93* cot94*...............*cot 178* cot 179 What is the answer? Plz describe with proper formulae

OpenStudy (experimentx):

write cot as cos/sin ... and , write all those terms as (90+x) on numerators and (180-x) on denominators and see what you get.

OpenStudy (anonymous):

\[\frac{ \cos(90+1).\cos(90+2).......\cos(90+88). \cos(90+89 }{ \sin(180-89). \sin(180-88).....\sin(180-2).\sin(180-1) }\] whats aftr this?? @experimentX

OpenStudy (experimentx):

yep ... what is value of \( cos(90+x) \) in terms of sin and what is value of \( \sin(180-x) \) in term of sin ??

OpenStudy (anonymous):

dont know...!! @experimentX

OpenStudy (experimentx):

cos(90+x) = -sin(x) sin(190-x) = sin(x)

OpenStudy (experimentx):

you know cos(A+B) and sin(A+B) formula?

OpenStudy (anonymous):

yes @experimentX

OpenStudy (experimentx):

yep ... use it to show cos(90+x) = -sin(x) sin(190-x) = sin(x) and use it on your problem.

OpenStudy (anonymous):

is the answer -1 @experimentX

OpenStudy (experimentx):

yep

OpenStudy (anonymous):

thnQ @experimentX

OpenStudy (experimentx):

yw